OpenStudy (anonymous):

Two forces with magnitudes of 25 and 30 pounds act on an object at angles of 10° and 100° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

@satellite73 @ganeshie8 @myininaya @jim_thompson5910

OpenStudy (dumbcow):

the resultant force is the sum of each of the components: \[R = <25\cos 10 + 30 \cos 100, 25\sin 10 +30\sin 100>\] magnitude: \[|R| = \sqrt{x^2 +y^2}\] direction: \[\tan \theta = \frac{y}{x}\]

OpenStudy (anonymous):

but how do you know exactly which terms were which? Thats kinda what im more confused on than anything lol

OpenStudy (dumbcow):

cos deals with horizontal force sin deals with vertical force

OpenStudy (anonymous):

okay so would i find the cos of 10 and multiply by 25... then do the same for "30cos100" then add them for the x variable? or would i do cos/sin^-1

OpenStudy (anonymous):

im not quite sure how to do this :/

OpenStudy (dumbcow):

yes, first option vector notation is <x,y>

OpenStudy (anonymous):

okay so thats all i would do? doesnt that come out to be a decimal though?

OpenStudy (dumbcow):

yes it says to round to 2 decimals

OpenStudy (anonymous):

ooooo I didnt know that was referring to the variables... so I should get.... x= 19.41 y= 33.88 ?

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@AccessDenied @Ashleyisakitty